In recent years I have been involved in interdisciplinary research. In this talk I will give a brief overview of some of these applied mathematics and interdisciplinary research projects.

Mathematics is the language of science, and as such, it can be used to describe and study various phenomena in diverse areas of science and technology. These problems in science and technology are a gold mine of mathematical problems.